OpenStudy (anonymous):

3. Solve the compound inequality. 3x + 12 ≥ –9 and 9x – 3 ≤ 33

OpenStudy (marigirl):

would u still like to know? or did u solve it

OpenStudy (triciaal):

multiply eqn 1 by 3 and rearrange to get 9x > or = -27 + 36 x>1 eqn 2 9x < or = 33 + 3 = 36 x < 4